OpenStudy (anonymous):

x-y=9 x+y=11 Determine if the solution is one, none, or infinitely many. If it's one then name the solution.

OpenStudy (trantom):

This is a set of paralell equations. You would use substitution for this. Solve for x on the first equation. x-y=9 Add y to both sides x=y+9 Now plug in y+9 for each x in the second equation. y+9+y=11 Combine like terms 2y+9=11 Subtract 9 2y=2 Divide by 2 y=1 Now go back to the first equation and plug in 1 for y. x-1=9 Add 1 x=10 You now know that x=10 and y=1. One solution. Hope this helped! :D

OpenStudy (anonymous):

Thank you! :)

OpenStudy (trantom):

Thats what openstudy is for :)